OpenStudy (anonymous):

in a series 2,3,5,6,7,11......there is no cube or square of any number. Find the 1000 digit in the series?

OpenStudy (anonymous):

where is 10?

OpenStudy (anonymous):

why 10?

jhonyy9 (jhonyy9):

but if you write 4 this will be one square of 2 and if will write 9 this will be one square of 3 or 16 will be one square of 4 or ... or 27 will be one cube of 3 but before if you write 8 this is one cube of number 2

OpenStudy (anonymous):

yes u r right sorry the actual series is as 2,3,5,7,10,11............

OpenStudy (anonymous):

i think its answer is 1039

jhonyy9 (jhonyy9):

sorry i think this is very difficult why you think that 1039 is the answer ?

OpenStudy (anonymous):

remove all the square from 2 to 33 & also remove cube of 1 to 10 number

jhonyy9 (jhonyy9):

ok and ???

OpenStudy (anonymous):

one minute

OpenStudy (anonymous):

taking square of 2^2=4 3^2=9 4^2=16 5^2=25 ..... .... 32^2=1024 as we know that 30^2 is 900 so we only calculate 31^2=961 which is less than 1000 & 32^2=1024 which is greater than 1000 so we taking only 31^2=961 so as this we know that there is we removing 2 to 31 numbers squares i.e. 30 from 2 to 1000 again we can caluculate cubes 2^3=8 3^3=9 4^3=16 5^3=125 6^3=216 7^3=343 8^3=512 9^3=729 10^3=1000 we not taking more than 10 cube total removes of cubes=9 then we can see that 64 & 729 is as two numbers which is comman to both so we can total remove=30+7=37 so we remove 999-37=962 then we know that 1001 is the 962th term then we want to find 1000th term so we add 1001+38 to find 1000th term i.e. 1039 but in this there is one more square term is to be there which want to be remove that is 32^2 which is 1024. so definetly the 1000th term is 1040 hence the answer is 1040

jhonyy9 (jhonyy9):

4 cubed not is 16 is 64

OpenStudy (anonymous):

oh sorry by mistak but answer is right

jhonyy9 (jhonyy9):

i like belive it but i am not sure in - sorry

OpenStudy (anonymous):

so u find answer